OpenStudy (anonymous):

If the difference of the roots of the equation x^2+px+q=0 be unity , then (p^2+4q^2) equals to

OpenStudy (anonymous):

(a) (1+2q)^2 (b)1-2q)^2 (c)4(p-q)^2 (d)2(p-q)^2

OpenStudy (vishweshshrimali5):

\[\large{\alpha + \beta = -p}\tag{1}\] \[\large{\alpha \beta = q}\tag{2}\] \[\large{\alpha - \beta = 1}\tag{3}\]

OpenStudy (vishweshshrimali5):

Here I am assuming that the roots are alpha and beta

OpenStudy (vishweshshrimali5):

Now using equations 1, 2 and 3: \[\large{(\alpha + \beta)^2 = (\alpha-\beta)^2 + 4\alpha\beta}\] \[\large{\implies p^2 = 1 + 4q}\]

OpenStudy (vishweshshrimali5):

Well appears the this equation is not useful for now ;)

ganeshie8 (ganeshie8):

option a satisfies above equation :P

OpenStudy (vishweshshrimali5):

^^^ Perfect method :P

ganeshie8 (ganeshie8):

weird way to cook up options lol

OpenStudy (vishweshshrimali5):

Lol :D

OpenStudy (vishweshshrimali5):

(a) (1+2q)^2 = 1 + 4q^2 + 4q = p^ + 4q^2

OpenStudy (vishweshshrimali5):

GOT IT !!!! :P

OpenStudy (anonymous):

wait a min

OpenStudy (vishweshshrimali5):

Solve the option a and use this equation: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 Now using equations 1, 2 and 3: \[\large{(\alpha + \beta)^2 = (\alpha-\beta)^2 + 4\alpha\beta}\] \[\large{\implies p^2 = 1 + 4q}\] \(\color{blue}{\text{End of Quote}}\) You will get the answer

OpenStudy (anonymous):

yes got

OpenStudy (vishweshshrimali5):

Great :D